OpenStudy (anonymous):

how do you convert 8.41 g/ml to kg/l "Please Help"!!!!!

OpenStudy (***[isuru]***):

Hi, first of all u need to be aware of the fact that milli(m) = 1/10^3 kilo(k) = 10^3 so , 8.41 * g/ml = 8.41* g/( 1/10^3)*l to convert g to kg we have g = 10^3 * (1/10^3) * g = (1/10^3) *kg therefor 8.41 * g/ml = [ 8.41* (1/10^3) *kg ]/ ( 1/10^3)*l = 8.41 *kg/l

OpenStudy (anonymous):

tank u:)

OpenStudy (***[isuru]***):

u r welcome!